OpenStudy (anonymous):

Admission to the zoo is $6 for children and $9 for adults. On a certain day, 1480 people go to the zoo and $10,500 is collected. How many children and how many adults went to the zoo?

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

I think children:940 Ithink adult :540

OpenStudy (anonymous):

i should make a template for this question lets cut right to the chase solve \[9x+6(1480-x)=10500\] and that will give you the number of adutls

OpenStudy (anonymous):

adult is 540 right

OpenStudy (anonymous):

which does make the number of children 940 so you got it

OpenStudy (mathmate):

This type of problem can often be solved mentally, as follows: Assume all tickets are for children, that brings in 6(1480)=9000-120=8880. We're missing 10500-8800=1620 in revenue. Changing each child to adult brings in $3, so need 1620/3=540 adults, so 1480-540=940 children.

OpenStudy (anonymous):

thxs to all closing the question in one min. and can yall help with one more question